# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

### v4

@@ 2.3.4 Accuracy of Determinstic Forecasts
where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
@@ 2.3.5 Accuracy of Probabilistic Forecasts
where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm. The second term in the equation is the mean pairwise distance between the scenarios.
3 Experimentation and Results
Compared to theoretical analysis, practical setups are more complex. They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization. This complexity is due to the necessity for cooperative behavior across multiple buildings. Additionally, various temporal dependencies emerge from switching costs and system feedback. To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
According to the decision matrix in Figure 1, at every time step the forecast and optimization plans can be updated or retained. Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$, respectively. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment. The result is shown in Tables 2(a) and 2(b). The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
(a) deterministic FHC algorithm with point forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
@@ 3 Experimentation and Results
(b) stochastic FHC algorithm with probabilistic forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
@@
Table 2: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.4. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours. For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [27]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period. This observation is notable as the most expected behavior is that best performance is achieved with shortest revision periods for both forecast types, as more up-to-date forecasts are provided.
These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
@@
(a) deterministic FHC algorithm
with point forecast
Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
$cost(FHC(v))$ 0.95 -0.94 -0.27
$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
(b) stochastic FHC algorithm
with probabilistic forecast
Correlation Coef. $ES$ $MASC_{V}$ $MASC_{H}$
$cost(FHC_{stoch.}(v))$ 0.63 -0.86 -0.29
$cost(FHC_{stoch.}(v))$ w/ switching costs 0.71 -0.06 0.47
Table 3: Correlation Coefficients for Forecast Metrics and Optimization Score
4 Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis in A.3, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@ 4 Discussion
The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions. In Figure 6, we can see that the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot. The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue). The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals. Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes. We assume that the oracle information given optimizing agent the confidence to charge at maximum capacity when the price and emissions are low. However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
@@
Figure 6: The difference in net load between FHC with and without switching costs. Actual net load is shown in yellow. The base load is shown in green. The net load following the perfect forecast is shown in red. The electricity price is shown in blue.
Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction. This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run. In line with the theoretical analysis in Appendix A.3, the upper bound on the competitive difference of the FHC algorithm is dependent on the trade-off between switching costs and forecast errors.
While the MPC objective functions factors in the switching costs, we identify two factors that can lead to underperformance of FHC with low commitment period ($v_{O}=1$), particularly when probabilistic forecasts are used. Firstly, when the policy is computed for a finite period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the previously executed decision may be suboptimal given the updated information. A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable. Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier. The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances. However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
Other important observation is that the FHC algorithm with longer commitment periods only becomes viable when the optimization is computed for a set of scenarios, as we can see from the results in Figure 5. It is hypothesized, that the aggregation of policies over different scenarios provides a more robust policy that hedges against the forecast errors. Generally, within a forecast horizon, the nearest steps are predicted with higher accuracy than further steps. Longer commitment periods aggravate the impact of forecast errors due to time coupling, as the suboptimal decisions made in the past are amplified in the future. A stochastic FHC algorithm is better equipped to handle errors in the forecast, as it aggregates policies over different scenarios, therefore, the impact of forecast errors is mitigated and longer commitment periods are more viable.
@@ 5 Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. Our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs. Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
Looking ahead, there are questions remaining to be addressed. For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy. Further research into the link between forecast quality and optimization performance is needed. Parameters such as forecast horizon, commitment period, and the number of scenarios are subject to a study for deeper understanding of their impact on the optimization performance and dependencies to the forecast accuracy and stability. It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading. From literature and current findings, methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors, however, they require more computational resources and are more complex to implement. An interesting avenue for future research would be defining the criteria for problems where these methods are most effective. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
@@ A.1 Optimization score
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (10)
@@
$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
@@
$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (16)
@@ A.2 Optimization Formulation
$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (19)
@@
  • 1.
    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • 2.
@@
A.3 Theoretical Analysis of Performance Bounds
In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$. We focus on the cost difference between the optimal policy and the policy produced by the FHC algorithm. Specifically, we examine the impact of the commitment $v$, which defines how often the policy and forecast are updated. Our goal is to show how the performance bound is influenced by a trade-off between forecast accuracy and switching costs.
To illustrate this, we consider two models of forecast error: independent and identically distributed (i.i.d.) noise, noise with exponentially decaying correlation and a probabilistic forecast. We start with the deterministic case (point forecast) and then extend to the probabilistic case (set of scenarios). Following the derivation in Chen et al. [18], we derive the average-case performance of the FHC algorithm by taking the expectation of the cost difference with respect to the prediction noise $(e(t))^{T}_{t=1}$.
Here, OPT represents the optimal policy that would achieve the lowest possible cost if there were perfect information and no prediction errors.
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (26)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition. The two noise models (i.i.d. and exponentially decaying correlation) are explained below.
Independent and Identically Distributed Noise
The i.i.d. noise is a simple and understandable baseline for forecast error, without time dependency. For i.i.d. noise, we have:
$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
The statistical properties of prediction errors in this context are:
$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
$\operatorname{trace}(R_{e})=\sigma^{2}$
This leads to the upper bound on the FHC cost:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\sigma^{\alpha}$ (28)
The upper bound on the competitive difference depends on $v$ only in the first term (in the denominator). Therefore, the optimal cost is minimized when $v=H$, i.e., the commitment level reaches the maximum allowable duration and corresponds to the forecast horizon length. The first term represents switching costs, so a shorter commitment period leads to more frequent updates and higher switching costs. The second term, representing i.i.d. noise, is independent of $v$.
Exponentially Decaying Correlation
A more realistic model of prediction noise uses correlations that decay exponentially over time. For this, given $a<1$, the Frobenius norm of the correlation structure is:
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (29)
We then express the prediction error covariance as:
$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\left\langle R_{e}^{1/2},f(s)\right\rangle^{2}$
$\leq\sum_{s=0}^{v}\left(\left\|R_{e}^{1/2}\right\|_{F}\|f(s)\|_{F}\right)^{2}$
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
Given that $h$ exhibits G-Lipschitz continuity ($\alpha=1$), the norm of the correlation function $\|f_{v}\|$ is bounded by:
$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (30)
Plugging this into the expression for the expected competitive difference of the FHC algorithm, we get:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (31)
$\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GT\sigma^{\alpha}\frac{c(1-a^{2(v+1)})/2}{1-a^{2}}$
$\leq\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GTc\sigma\left(\frac{a^{2v+2}-2}{2(a^{2}-1)}\right)$
To simplify, we group the parameters independent of $v$:
$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
Thus, the expected competitive difference of the FHC algorithm is:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{A}{v}+B(Ca^{2v}-1)$ (32)
The competitive difference behavior depends on two main components. The term $\frac{A}{v}$ decreases with $v$, while the term $B(Ca^{2v}-1)$ increases with $v$. Since $Ca^{2v}$ approaches 0 as $v$ increases, the second term is bound by $B$. Therefore, the minimal competitive difference is a trade-off between switching costs (first term) and forecast errors (second term).
Stochastic Optimization
For stochastic optimization with Fixed Horizon Control (FHC) applied across $n$ scenarios, we examine how the stochastic nature influences the competitive difference compared to an optimal solution. The expected cost for the stochastic algorithm across all scenarios is:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))=\sum_{i=1}^{n}\mathbb{E}\operatorname{cost}(FHC(v,i))$ (33)
Taking the expectation of the competitive difference across all scenarios:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (34)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\left[\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}\right]$
In the stochastic case, the forecast errors across the scenario ensemble would be lower in expectation than in any singular, deterministic scenario. This is because the stochastic approach averages out the errors across multiple scenarios, reducing the overall forecast error. Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
A.4 In-sample Performance
(a) Performance of the FHC algorithm with point forecast.
Study over combinations of forecast and optimization commitment between 1 and 12
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12

### v5

@@ 4.3.3 Accuracy of Probabilistic Forecasts
4.4 Results
According to the decision matrix in Table 1, at every time step the forecast and optimization plans can be updated or retained. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment, $v_{F}$ and $v_{O}$, respectively. The result is shown in Table 3. The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
Table 3: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours
Optimization Commitment Period ($v_{O}$)
Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
Panel A: Deterministic FHC with point forecast
@@ 4.4 Results
Panel B: Stochastic FHC with probabilistic forecast
@@
Note: Diagonal elements represent equal commitment periods for both forecast and optimization. Lower values indicate better performance. Empty cells (—) represent invalid combinations where $v_{O}>v_{F}$.
@@
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 6, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in A.3. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours. For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in Nweye et al. [2022]. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation. Nevertheless, as seen in the optimization score plots, the global minimum is reached at a shorter commitment period, indicating that updating the schedule more frequently is not only beneficial for using the most accurate forecast but also for correcting the previous decisions.
Furthermore, it is observed that the presence of switching costs has a significant impact on the relation of performance to the commitment period. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. Similar findings are devised from the average-case analysis in Section 3.1, where higher switching costs can potentially lead to a formation of a minimum beyond lowest commitment period. The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period. This observation is notable as, conventionally, the best performance is achieved with the shortest revision periods. This observation is also consistent with the theoretical analysis in Section 3.2, as the stochastic optimization reduces sensitivity to forecast errors, leading to a more stable performance.
These observations are supported by the correlation coefficients in the Table 4 with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
@@
Figure 7 demonstrates an example from the Citylearn case when the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot. The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue). The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals. Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes. We assume that the oracle information gives the optimizing agent the confidence to charge at maximum capacity when the price and emissions are low. However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
@@
Table 4: Correlation Coefficients between Forecast Metrics and Optimization Performance
Point Forecast (Deterministic FHC) Probabilistic Forecast (Stochastic FHC)
Correlation Coefficient $MAE$ $MAC_{V}$ $MAC_{H}$ $ES$ $SDC_{V}$ $SDC_{H}$
Without switching costs 0.95 -0.94 -0.27 0.63 -0.86 -0.29
With switching costs 0.70 -0.62 0.17 0.71 -0.06 0.47
@@
Figure 7: Net load comparison: Without switching costs, energy management with perfect forecast (yellow) generates high peaks versus conservative forecast-based control (blue). With switching costs, the policy smoothes fluctuations. The base load, being the prediction target variable, is shown in blue. Price and emissions signals are shown in the bottom subplot.
5 Discussion
This study provides theoretical and empirical evidence for the nuanced role that commitment periods play in forecast-based optimization under switching costs. Our results validate the theoretical performance bounds developed in Section 3, showing that the trade-off between forecast inaccuracy and switching costs emerges most prominently when switching penalties are non-negligible. Notably, this trade-off is not universal: in problems without meaningful switching costs, shorter commitment periods are almost always preferred. However, when switching costs are significant, such as in systems with physical constraints, ramping penalties, or contract-driven operational inertia, the choice of commitment level becomes a key decision variable in its own right.
The presence of switching costs changes the optimization landscape in fundamental ways. It introduces temporal coupling between successive decisions, requiring current actions to anticipate the cost of future revisions. This often leads to smoother and more conservative policies, as illustrated in Figure 7, the addition of switching costs results in a visibly flattened net load curve. Such behavior is desirable in many real-world settings: battery degradation, generator ramping, thermal comfort zones, and load shifting incentives all impose implicit or explicit switching penalties that must be carefully managed. Although such costs are context-specific and often application-dependent, they are far from rare in modern energy systems.
Importantly, we show that this trade-off is more tractable under stochastic optimization. Section 3.2 demonstrates that scenario averaging introduces a form of regularization, lowering the effective sensitivity to forecast errors. This effect is reflected in both the theoretical bounds and the experimental results. In practice, this means that higher commitment levels, normally problematic due to the accumulation of outdated forecast errors, can be made viable when robust scenario-based policies are used. The performance gains from longer commitment periods become visible only when switching costs are high and stochastic optimization is employed, as shown in Figure 6. This aligns with the theoretical prediction that the second term in the cost bound (forecast error contribution) becomes less dominant when multiple forecast scenarios are aggregated.
From an operational perspective, this has several implications. First, minimizing switching costs is not solely about reducing the number of policy changes. It also requires designing forecasts that are stable across time. As formalized in Section 3.3, vertical forecast stability directly mitigates the switching cost penalty. Stability metrics such as Mean Absolute Change (MAC) or Scenario Distribution Change (SDC) can thus serve as practical proxies to monitor and optimize for reduced control variability in real systems. Second, while the importance of forecast accuracy has long been emphasized in predictive control, this study shows that stability is an equally critical, yet underappreciated property, especially when switching costs are present. Systems that update policies too frequently may suffer from over-correction, chasing short-term forecast noise at the expense of long-term efficiency.
@@ 6 Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. Our analysis shows that in systems where forecasts guide decision-making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts. Our findings have significant implications for the design of energy management systems. In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics. The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs. Furthermore, we propose $SDC$, a novel metric to evaluate the stability of scenario sets.
Our study has several important limitations that should be acknowledged. First, our theoretical analysis assumes convex cost functions and exponentially decaying forecast errors, which may not fully capture the complexity of real-world energy systems. Second, while we demonstrate the relationship between forecast stability and optimization performance on a single case study, further validation across different energy management contexts and geographical locations would strengthen our findings. Additionally, our definition of switching costs primarily focuses on grid ramping, which serves as a proxy for the more complex switching costs in real energy systems, including wear-and-tear, efficiency losses, and network fees. Future work should address these limitations by testing across diverse datasets, and exploring more sophisticated methods for enhancing forecast stability without compromising accuracy. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape. It would also be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
@@ A.1 Optimization score
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}}\right)$ (29)
@@
$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
@@
$D=\operatorname{avg}\left(\frac{R_{\text{ALG}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{ALG}}}{1-L_{\text{no\ battery}}}\right)$ (35)
@@ A.2 Optimization Formulation
$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{S}_{tt^{\prime}})$ (38)
@@
  • $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the look-ahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
@@
A.3 In-sample Performance
Table 5: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours (with data from in-sample buildings)
Optimization Commitment Period ($v_{O}$)
Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
Panel A: Deterministic FHC with point forecast